Using Wilson's and Fermat's Little Theorem
What is the remainder when:
1.) 3^{31}\div73 31 ÷ 7
2.) 5^{119}\div595 119 ÷ 59
3.) 2^{70}+3^{30}\div132 70 + 3 30 ÷ 13
4.) 53!\div6153 ! ÷ 61
5.) 149!\div139149 ! ÷ 139
Find all integers x such that: x^{86}\equiv6\left(\mod29\right)
1
Expert's answer
2020-09-29T16:42:03-0400

1) By Fermat’s Little Theorem,

"3^{7-1}\equiv1(\bmod7)"

Then

"3^{12}\equiv1(\bmod7)," "3^{18}\equiv1(\bmod7)," "3^{24}\equiv1(\bmod7)," "3^{30}\equiv1(\bmod7)"

"3^{1}\equiv3(\bmod7)"

"3^{31}\equiv3(\bmod7)"


2) By Fermat’s Little Theorem,

"5^{59-1}\equiv1(\bmod59)"

Then "5^{116}\equiv1(\bmod59),"

"5^{3}\equiv7(\bmod59)"

"5^{119}\equiv7(\bmod59)"


3) By Fermat’s Little Theorem,

"2^{13-1}\equiv1(\bmod13), 3^{13-1}\equiv1(\bmod13)"

"2^{60}\equiv1(\bmod13), 3^{24}\equiv1(\bmod13)"

Then

"2^{70}+3^{30}\equiv2^{10}+3^{6}\equiv1024+729\equiv"

"\equiv114+79\equiv193\equiv11(\bmod13)"

"2^{70}+3^{30}\equiv11(\bmod13)"


4) By Wilson's Theorem,

"(61-1)!\equiv-1(\bmod61)"

"(-1)(-2)(-3)(-4)(-5)(-6)(-7)53!\equiv-1(\bmod61)"

"53!(2)(3)(4)(5)(6)(7)\equiv1(\bmod61)"

"(2)(5)(6)=60"

"53!(-1)(3)(4)(7)\equiv1(\bmod61)"

"53!(3)(4)(7)\equiv-1(\bmod61)"

"(3)(4)(7)=84"

"53!(23)\equiv-1(\bmod61)"

Let's do Euclidean algorithm to compute "23^{-1}" "\bmod61"

"61=2(23)+15"

"23=15+8"

"15=8+7"

"8=7+1"

Hence

"1=8-7=8-(15-8)=2(8)-15=2(23-15)-15="

"=2(23)-3(15)=2(23)-3(61-2(23))="

"=8(23)-3(61)"

"23^{-1}\equiv8(\bmod61)"


Hence

"53!\equiv-8(\bmod61)"

"53!\equiv53(\bmod61)"


5) 139 is the  prime number. By Wilson's Theorem,

"(139-1)!\equiv-1(\bmod139)"

"(10)(9)(8)(7)(6)(5)(4)(3)(2)(1)(-1)\equiv"

"\equiv(9)(8)(6)(5)(4)(3)(-1)\equiv(9)(8)(3)(-19)(-1)\equiv"

"\equiv(27)(13)\equiv73(\bmod139)"

"149!\equiv73(\bmod139)"


6) By Fermat’s Little Theorem,

"x^{29-1}\equiv1(\bmod29)"

"86=3(28)+2"

"x^{86}\equiv x^2(\bmod29)"

Then

"x^{2}\equiv 6(\bmod29)"

"x^{2}\equiv 64(\bmod29)"

"x^{2}-64\equiv 0(\bmod29)"

"(x-8)(x+8)\equiv (\bmod29)"


"x\equiv 8(\bmod29)" or "x\equiv 21(\bmod29)"